Find the general solution of the following differential equations using method of undetermined coefficients: (ii) y''−2y'+y =20e^xsinx,


1
Expert's answer
2022-03-02T09:31:17-0500

Corresponding homogeneous differential equation


"y''\u22122y'+y =0"

Characteristic (auxiliary) equation


"r^2-2r+1=0"

"r_1=r_2=1"

The general solution of the homogeneous differential equation is


"y_h=c_1e^x+c_2xe^x"

Find the particular solution of the non homogeneous differential equation


"y_p=Ae^{x}\cos x+Be^x\sin x"

"y_p'=Ae^x\cos x-Ae^x\sin x+Be^x\sin x+Be^x\cos x"

"y_p''=Ae^{x}\cos x-2Ae^x\sin x-Ae^{x}\cos x"

"+Be^x\sin x+2Be^x\cos x-Be^x\sin x"


Substitute


"-2Ae^x\sin x+2Be^x\cos x-2Ae^x\cos x"

"+2Ae^x\sin x-2Be^x\sin x-2Be^x\cos x"

"+Ae^{x}\cos x+Be^x\sin x=20e^x\sin x"

"A=0"

"B=-20"

The general solution of the non homogeneous differential equation is


"y=c_1e^x+c_2xe^x-20e^x\sin x"